Uk's real GDP was 1 360 trillion pounds (£) in 2009 and 1 434 trillion pounds in 2010. Uk's population was 191.5 million people in 2009 and 193.3 million in 2010.
Calculate:
1. The economic growth rate .(3 marks)
2. The growth rate of real GDP per person .(5 marks )
3. The approximate number of years it takes for real GDP per person in the Uk to double if the 2010 economic growth rate and population growth rate are maintained. (4 marks)
1
Expert's answer
2019-08-05T12:01:17-0400

a. "\frac{1434-1360}{1360}"  =0.05441


b.The growth rate of real GDP per person="\frac{7.419-7.102}{7.102}"​ = 0.0446


c. The GDP per person = 4.46% it will double= using the rule of 72 = "\frac{72}{4.46}" = 16 year